OpenStudy (anonymous):

The graph below represents which system of inequalities? y less than or greater to -2x + 3 y less than or greater to x + 3 y greater than or equal to -2x + 3 y greater than or equal to x + 3 y less than or greater to -3x + 2 y less than or greater to -x + 2 y > -2x + 3 y > x + 3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

let me post the graph @AlexandervonHumboldt2

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

im assuming the green area is the shaded solution?

OpenStudy (amistre64):

define the eqautons of the 2 boundary lines

OpenStudy (amistre64):

and in each case, the shading is below(less than) the boundary line

OpenStudy (anonymous):

@iGreen help please~

OpenStudy (anonymous):

@epoweritheta will medal

OpenStudy (anonymous):

please help @Ilovecake @texaschic101

OpenStudy (texaschic101):

lets check out the answer choices.... y < = -2x + 3 set aside the inequality sign for now y = -2x + 3 slope = -2 y int = 3 x int ---- sub in 0 for y 0 = -2x + 3 2x = 3 x = 3/2 or 1 1/2 and we know it is shaded below the line.....so this is one of the lines on your graph. y <= x + 3 slope = 1 y int = 3 x int ....sub in 0 for y 0 = x + 3 x = -3 shaded below the line means it is less then here is your other line

OpenStudy (anonymous):

so is it a @texaschic101

OpenStudy (texaschic101):

I guess.....I see no A's.....it's the first 2 answer choices...or the first set of lines.... answer is : y <= -2x + 3 and y <= x + 3

OpenStudy (anonymous):

this is a @texaschic101 y less than or greater to -2x + 3 y less than or greater to x + 3

OpenStudy (anonymous):

this is b @texaschic101 y greater than or equal to -2x + 3 y greater than or equal to x + 3

OpenStudy (anonymous):

so which one do you think it is?:) @texaschic101

OpenStudy (anonymous):

did you mean -2 ? @texaschic101 and the sign < with a line under it?

OpenStudy (texaschic101):

the sign < with a line under it......yes, I can't do that on my computer so I just put the equal sign beside it. answer is : y (less then or greater) then -2x + 3 and y (less then or greater) then x + 3

OpenStudy (anonymous):

awesome will fan u help with 1 more pls? @texaschic101

OpenStudy (texaschic101):

ask away

OpenStudy (texaschic101):

sorry about the wait....got a little distracted